The length of a rectangle is 2 feet less than twice its width. If the area of the rectangle is 100 feet, what is the width? Type in the larger of your two answers, to the nearest hundredth.
Answer by fractalier(6550)

Area A = LW = 100. We also have
L = 2W - 2
Substituting we get
(2W - 2)W = 100
2W^2 - 2W = 100
W^2 - W - 50 = 0
Using the quadratic formula we find that
W = 7.59 feet